OpenStudy (anonymous):

What is 98% more than 1170?

OpenStudy (faiqraees):

find the 98 percent of 1170 and add that into 1170? Clear or more explanation?

OpenStudy (wcrmelissa2001):

Lemme justexplain in detail to make sure... Ok so first you need to find 1%. Since 1170 is 100%, to find 1% 1170/100 = 11.7 98% more than 100% is 198%. 11.7*198=2,316.6

OpenStudy (anonymous):

That's exactly what I thought! I was wondering if I had to multiply 1170 and 98 to get 98% of 1170 and then add the total to 1170? It just didn't sound too right so I wasn't sure. So 98% of 1170 is: 98 x 170 = 114660 114660 + 1170 = 115830 Is that correct?

OpenStudy (anonymous):

@wcrmelissa2001 answer is a bit confusing... I mean I understand, but how do you apply that to the question?

OpenStudy (faiqraees):

no you have to find out 98 PERCENT of 1170 which is \[\large\rm \frac{98}{100}*1170=1146.6 \]Now add this into 1170 \[\large\rm 1170+1146.6=2316.6 \]

OpenStudy (faiqraees):

Got it? Or more explanation?

OpenStudy (abdullahm):

Another way to do it: 98% more than 1170 That's the same as calculating: (98% * 1170) +1170 Which is the same as: 198% * 1170 That's because 98% more of 1170 is 198% of 1170.

OpenStudy (phi):

***That's exactly what I thought! I was wondering if I had to multiply 1170 and 98 *** Just to repeat what people have said: you don't multiply by 98 , you multiply by 98/100 98 per cent is Latin for 98 per hundred or 98 out of a hundred. i.e. 98/100

OpenStudy (phi):

also, the question What is 98% more than 1170? is "leaving out" some words (which they assume you can figure out) To be more clear, they could say what is 98% (of 1170) more than 1170?

OpenStudy (sparrow2):

actually it depends how you ask a question. suppose that number is x and it is 98%(of this numebr,not 1170) more than 1170 so we got x-x*0,98=1170 x=1170*50

OpenStudy (sparrow2):

so let's consider the following example: suppose we have two numbers 8 and 10.if you say,how many persent is 8 less then 10? we know that the difference is 2,but when we asked for percent what should we take 2/8 =20% or 2/10=25%? i think the first one,because you speak form 8's perpective.if i sad with how many percent is 10 greater then 8, i would choose 2/10=25% because we speak from 10's percpective.

OpenStudy (sparrow2):

for your example you speak from the perspective of that unknown number

OpenStudy (anonymous):

Ohhhh, okay! Thank you guys! I understand now! :)